OpenStudy (anonymous):

You toss 3 fair coins in the air at once. a) List the sample space.

OpenStudy (anonymous):

would it be HT, HT, HT

OpenStudy (anonymous):

no

OpenStudy (anonymous):

there are two possibilities for each coin and by the counting principle there are therefore \(2\times 2\times 2=2^3=8\) possible outcomes

OpenStudy (anonymous):

since 8 is not such a large number, we can list all of them for example, you could get h , h, h

OpenStudy (anonymous):

the whole list looks like h, h, h h, h, t h, t, h t, h, h h, t, t t, h, t t, t, h t, t, t

OpenStudy (anonymous):

you can think of it as tossing a penny, a nickel and a dime, although it does not matter but if you think of it that way, then for example the first one would be the penny, the second the nickel and the third the dime in this reckoning h, t, t for example means you got a head on the penny, a tail on the nickel and a tail on the dime

OpenStudy (anonymous):

thank you.